OpenStudy (ledah):

Finish writing the equation of the line , given the points (6, -5) and (-2, 7).

OpenStudy (ledah):

Sorry its a bit hard to see :c

OpenStudy (jit4won):

m = ( change in y coordinates )/ (change in coordinates)

OpenStudy (jit4won):

m = (-5 - 7) / ( 6 +2 ) = - 3/ 2

OpenStudy (jit4won):

use m( x- x0) = y-y0

OpenStudy (jit4won):

x0 = 6 y0 = -5

OpenStudy (jit4won):

solve yourself

OpenStudy (ledah):

thanks